OpenStudy (anonymous):

Which polynomial is a perfect square trinomial? 36x2 − 18x − 9 4a2 − 28a − 49 25b2 − 40b + 16 4x2 − 6x + 9

OpenStudy (magbak):

A

OpenStudy (magbak):

No it is D

OpenStudy (anonymous):

LOL xD

OpenStudy (anonymous):

both wrong..

OpenStudy (magbak):

Trust me it is D

OpenStudy (anonymous):

Hahaha :D

OpenStudy (anonymous):

It's C xD

OpenStudy (anonymous):

\[(ax+b)^2 = ax^2 +2ab + b^2\]

OpenStudy (nincompoop):

why are we giving out answers?

OpenStudy (magbak):

6 divided by 2 it 3. And 3 squared is 9

OpenStudy (anonymous):

25b2 − 40b + 16 (5b)^2-2(5b)(4)+(4)^2 (5b-4)^2

OpenStudy (cwrw238):

its C

OpenStudy (anonymous):

Yaaaayyyy I won xD

OpenStudy (jhannybean):

Take b. \[\large 4a^2-28a-49\]\[\large a^2 =4a^2\]\[\large b^2 = 49\] \[\large a^2-b^2\] Doesnt fit the form. try c. \[\large a^2-b^2 = a^2 -2ab+b^2 \]\[\large a^2 = 25a^2 \ \therefore \ a= 5a\]\[\large b^2 = 16 \ \therefore \ b=4\]put it into our form of difference of squares.

OpenStudy (nincompoop):

listen to professor beanhead she got this

OpenStudy (anonymous):

I'll explain :) 36x2 − 18x − 9 The formula is (x-2)^2 = x^2-4x+4 So plus should be in the end but there is minus so this can't be the perfect square xD 4a2 − 28a − 49 Same case minus is there instead of plus 25b2 − 40b + 16 This is perfect square because 25b2 − 40b + 16 (5b)^2-2(5b)(4)+(4)^2 (5b-4)^2 4x2 − 6x + 9 This even cant be the perfect square So, option C is the answer xD

OpenStudy (jhannybean):

Good job.